OpenStudy (anonymous):

Solve by substitution. 3x − 7y = 35 x = −1 − 4y

hero (hero):

y = (35-3x)/-7 y = (x+1)/-4 y = y (substitute) -7(x+1) = -4(35-3x) -7x-7 = -140+12x 19x = 133 x = 133/19 x = 7 y = -2

hero (hero):

Any questions?

OpenStudy (anonymous):

not at this time, thank u